Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Proability

Expert replies
by hariharakarthi » Mon Aug 03, 2009 9:02 am
A certain deck of cards contains 2 blue cards, 2 red cards, 2 yellow cards, and 2 green cards. If two cards are randomly drawn from the deck, what is the probability that they will both are not blue?
Join the discussion
Source: — Problem Solving |

by gmat_dest » Mon Aug 03, 2009 9:11 am
Probability of blue card = 2/8

Probability of not blue card = 6/8

Probability of both cards being not blue = 6/8 * 5/7 = 15/28
Join the discussion

by tom4lax » Tue Aug 04, 2009 6:28 pm
Wouldn't the easiest way to approach this problem be to determine the probability that both are blue?

1st card blue = 2/8
2nd card blue = 1/7

2/8 * 1/7 = 1/28 both are blue

1 - 1/28 = 27/28 neither are blue

Whats the OA?
Join the discussion

by ankitns » Wed Aug 05, 2009 6:22 am
If the probability of both being blue is 1/28.

1 - 1/28 = 27/28 will give us the probability that BOTH OF THEM ARE NOT BLUE at the same time. This means that you could have 1 blue and 1 other color...or 2 other colors. So this approach will not work.


tom4lax wrote:Wouldn't the easiest way to approach this problem be to determine the probability that both are blue?

1st card blue = 2/8
2nd card blue = 1/7

2/8 * 1/7 = 1/28 both are blue

1 - 1/28 = 27/28 neither are blue

Whats the OA?
Attempt 1: 710, 92% (Q 42, 63%; V 44, 97%)
Attempt 2: Coming soon!
Join the discussion

by umaa » Wed Aug 05, 2009 6:27 am
Probability of both are blue = 2/8 * 1/7 = 1/28

So, the probability of both are not blue is, 1-(1/28) = 27/28
Join the discussion

by mohitsharda » Wed Aug 05, 2009 7:39 am
tom4lax wrote:Wouldn't the easiest way to approach this problem be to determine the probability that both are blue?

1st card blue = 2/8
2nd card blue = 1/7

2/8 * 1/7 = 1/28 both are blue

1 - 1/28 = 27/28 neither are blue

Whats the OA?
I think that the case that you are taking is that both are blue. and then subtracting it to get the case when both are not blue. But the question should be interpreted as not even a single card is blue. In this reasoning you are taking cases when 1 is blue and other is not and hence you are getting 27/28

gmat_dest is right in the way he has solved it.. according to me
MS
Join the discussion

Re: Proability

by ankitns » Wed Aug 05, 2009 7:41 am
Could you please clarify the last sentence "what is the probability that they will both are not blue?"

Do you mean that they are BOTH not blue

or

Do you mean that NEITHER of them are blue (which means that at most 1 of them can be blue).

Depending on that, the question changes and hence the probability changes.

Thanks.

quote="hariharakarthi"]A certain deck of cards contains 2 blue cards, 2 red cards, 2 yellow cards, and 2 green cards. If two cards are randomly drawn from the deck, what is the probability that they will both are not blue?[/quote]
Attempt 1: 710, 92% (Q 42, 63%; V 44, 97%)
Attempt 2: Coming soon!
Join the discussion

by tom4lax » Wed Aug 05, 2009 2:18 pm
Okay, got it guys, thanks.

Question is asking probability that NEITHER is blue.

2/8 first card is blue, so 6/8 it isnt

5/7 second card is not blue

5/7 * 6/8 = 15/28
Join the discussion